OpenStudy (anonymous):

Hello everyone=) I have couple questions from my homework. I will be happy if someone can help me with them: 1) The points on a scatterplot lie very close to the line whose equation is y = 2 − 5x. The correlation between x and y is close to A. 2 B. -5 C. -1 2)For a biology project, you measure the weight in grams and the tail length in millimeters of a group of mice. The correlation is r = 0.7. If you had measured tail length in centimeters instead of millimeters, what would be the correlation? (There are 10 millimeters in a centimeter.) A. 0.7/10 = 0.07 B. 0.7 C. (0.7)(10) = 7

OpenStudy (anonymous):

Ie to say if the scatterplot line was plotted, the gradient of that line would be similar to the y = 2 - 5x one. If the gradient of y = 2-5x is -5, thus gradient of scatterplot would be ~~ -5. Correlation will not change with units

OpenStudy (anonymous):

thank you!